Two cars leave Orlando heading towards Atlanta. The first car is going 50 mph and the second car is gping 70 mph. The first car leaves at 9am and the second car leaves at 11am. How long will it take for the second car to pass the first car? Thank you very much.
Answer by Edwin McCravy(20054)

Two cars leave Orlando heading towards Atlanta. The first car is going 50 mph and the second car is gping 70 mph. The first car leaves at 9am and the second car leaves at 11am. How long will it take for the second car to pass the first car? Thank you very much.

How to do it in your head:
When the second car leaves, the first car is d=rt=50×2=100 miles away.  The
second car's approach rate is 70-50 = 20 mph, and it will take him 
t=d/r=100÷20 or 5 hours.

How to do it by algebra:

Make this chart:

              distance     rate   time     
first car                  
second car               

Fill in the rates (speeds)


              distance     rate   time     
first car                   50      
second car                  70      

Let t = the time for the second car (which is what we want to find)
Then since the first car left 2 hours earlier, its time is t+2.
Fill those in


              distance     rate   time     
first car                   50     t+2  
second car                  70      t

Fill in the distances using d=rt

              distance     rate   time     
first car      50(t+2)      50     t+2  
second car       70t        70      t

When the second car catches the first car, the
distances are equal:
                 
                  50(t+2) = 70t
                50t + 100 = 70t
                     -20t = -100
                        t = 5 hours

Edwin